Q. A railway flat car,whose mass together with the artillery gun is M , moves at a speed V . The gun barrel makes an angle ? with the horizontal. A shell of mass 'm' leaves the barrel at a speed v,relative to the barrel.The what is the speed of the flat car in order to stop after the firing?

Grade:11

2 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Raj, 

Let the angle be 'theta'.  from conservation of horizontal component of linear momentum, we have

m(vcos(theta)+V)=MV, V= mvcos(theta)/(M-m)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras

Archit Uttarwar
5 years ago
 
Dear Raj, 
 
Let the angle be 'theta'. from conservation of linear momentum in horizontal direction  we have
m(vcos(theta)+V)=(M+m)  V
V= mvcos(theta)/(M)